OpenStudy (anonymous):

Please helppppp!!!!Mike and Larry are making four tables for the eight grade store. The rectangular top of each table will have a length of 3 3/4 feet and width of 2 3/5 feet.Find the perimeter of the four table tops combined.

OpenStudy (hockeychick23):

i would make these improper fractions first: 3 3/4= 15/4 2 3/5= 13/5 Now we need a common denominator which will be 20 in this case 15/4*5/5= 75/20 13/5*4/4= 52/20 75/20+75/20+52/20+52/20= 254/20= 12.7 or 12 7/10

OpenStudy (hockeychick23):

since there are 4 table tops we need to take 12.7+12.7+12.7+12.7

OpenStudy (anonymous):

so is that the perimeter of four table tops? 50.8 feet

OpenStudy (hockeychick23):

correct :)

OpenStudy (anonymous):

Each table has two 3.75 foot long sides and two 2.6 foot long sides. The perimeter of one table would be 2*3.75+2*2.6 = 7.5 + 5.2 = 12.7 feet per table. Times 4 tables = 12.7 x 4 = 50.8 feet